# 11footnotetext: Equal Contributions

### v2

@@ 4.3 Comparisons
  • Take-over-Correction (HG-DAgger) [21] is another common correction strategy where human demonstrators monitor policy execution and take complete control when failures are anticipated. We implement Take-over-Correction on our Compliant Intervention Interface by cleaning up command buffer to the compliance controller and switching stiffness to zero upon correction starts, so the robot policy does not affect the robot during correction. However, as explained in § 3.1, take-over correction introduces an abrupt transition around control authority switching, which may cause distributional discontinuities in the training data.
@@ 4.4 Key Findings
Finding 2: Residual policy allows additional useful modality to be added during correction. [Compliant Residual] policy performs significantly better than other methods without force (45% higher success rate than the best position-only baseline on the book task and 20% higher on the belt task) as it can both take in force feedback that indicates critical task information and predict adequate contact forces to apply. For example, the last stage of the book flipping task requires the robot to firmly push the book against the shelf wall to let it stand on its own. [Compliant Residual] policy predicts large pushing forces at this stage to make the books stand stably with a 100% success rate, while [Residual]’s success rate drops from 70% to 35% (§ A.1). The second stage of the belt assembly task (threading the belt on the large pulley) requires delicate belt height adjustments under ambiguous visual information due to occlusions and the lack of depth. [Compliant Residual] policy learns to move along the pulley to find the slot when the finger touches the top of the pulley.
@@
Finding 3: Smooth On-Policy Delta data enables stable residual policy. [Compliant Residual] policy trained with [On-Policy Delta] data has 45% and 30% higher success rates than [Compliant Residual] & [Take-over-Correction] on the two tasks respectively. Both residual policies trained with [Take-over-Correction] data sometimes exhibit large noisy motions that trigger task failures, such as retracting the fingers too early in the book flipping task, and pulling the belt forcefully above the large pulley in the belt assembly task. On the contrary, residual policies trained with [On-policy Delta] data have much smoother action trajectories and better reflect human’s correction intentions, providing suitable magnitudes of corrections.
@@
Finding 4: Retraining base policy is stable but learns correction behavior slowly. Retraining from scratch with the initial and correction data together leads to policies with stable motions. However, its behavior is less affected by the small amount of correction data compared to the dominant portion of initial data, leading to insignificant improvements over the base policy (1.67% success rate drop on the book task and 5% increase on the belt task, averaged across all data collection methods).
@@
Finding 5: Finetuning base policy is unstable. Policy finetuning with either correction data has the worst performance across all policy update methods and even underperforms the base policy (30% success rate drop on the book task and 15% drop on the belt task, averaged across all data collection methods). The finetuned policy predicts unstable and noisy motions, quickly leading to out-of-distribution states, such as inserting too high in the book flipping task and drifting away from the board in the belt assembly task. This is likely due to the distribution mismatch between the base policy training distribution and correction data distribution, causing training instabilities.
@@ 4.5 Ablations
Training frequency and batch size. One important parameter in DAgger is the batch size between policy updates. With a smaller batch size, the policy is updated more frequently, then new correction data can better reflect the most recent policy distribution. However, DAgger with small batch sizes is known to suffer from catastrophic forgetting [25, 10] since it finetunes neural networks on data with non-stationary distribution. Common solutions include retraining the residual policy at the end of DAgger using all available correction data collected from all the intermediate residual policies [37]. Another way is to rely on the base policy training data as a normalizer [15]. In this work, we empirically found that a much simpler method works the best: single-batch training, i.e., we train the residual only once. All correction data is obtained when only running the base policy. We compare our single-batch (batch size = 50) DAgger with a small batch size version, where we warm up the residual with 20 episodes of initial correction data, then update every ten more episodes for three times.
Figure 7: Effect of Training Frequency and Sample. Single-batch update (batch size=50) leads to more stable training and dense sampling after correction starts achieves better performance.
Finding: Single-batch DAgger is more suitable for training Compliant Residual Policy. The small-batch training becomes unstable and the demonstrator needs to provide large magnitudes of corrections as the number of iterations increases. During evaluation, the final policy always fails by inserting too high, while our single-batch policy achieves a 100% success rate with the same amount of data and training epochs.
Sampling strategy during training. The start of a human intervention contains critical information of the timing and direction of correction. Accurate delta action predictions right after correction starts are important for reactive corrective behaviors and staying in distribution. We investigate three strategies for sampling from online correction data during training: 1. Uniform sample, where the whole episode is sampled uniformly. 2. Denser sample around the start of a human intervention, and 3. denser sample only after the human intervention starts. For 2 and 3, we uniformly increase the sample frequency four times for a fixed period before and/or after intervention starts.
@@ 5 Conclusion and Discussion
In this work, we evaluate practical design choices for DAgger in real-world robot learning, and provide a system, CR-DAgger, to effectively collect human correction data with a Compliant Intervention Interface and improve the base policy with a Compliant Residual Policy. We demonstrate the effectiveness of our designs by comparing them with a variety of alternatives on two contact-rich manipulation tasks.
Limitations and Future Work. The base policy should have a reasonable success rate for the residual policy to learn effectively. From our experiments, we recommend starting to collect correction data for the residual policy when the base policy has at least $10\%\sim 20\%$ success rate. A future direction is to derive theoretical guidelines for the trade-off between the base and residual improvements.
Throughout this work, we use a MLP as the action head of our Compliant Residual Policy and directly regress the actions. Although it works well in our tasks, it may experience difficulty for tasks that involve more distinctive action multi-modalities. More expressive policy formulations, such as Flow Matching [26, 5] might be useful for these tasks.
@@ 6 Acknowledgment
We would like to thank Eric Cousineau, Huy Ha, and Benjamin Burchfiel for thoughtful discussions on the proposed method, thank Mandi Zhao, Maximillian Du, Mengda Xu, and all REALab members for their suggestions on the experiment setup and the manuscript. This work was supported in part by the NSF Award #2143601, #2037101, and #2132519, Sloan Fellowship, and Toyota Research Institute. We would like to thank Google and TRI for the UR5 robot hardware. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Stage-Wise Success Rate
Both the book flipping task and the belt assembly task can be further divided into three key stages. Fig. 8 below is a more detailed version of Fig. 6, which reports all the task success rates by stages.
@@ A.1 Stage-Wise Success Rate
A.2 Use of Base policy Action as Input to the Residual Policy
One design choice for the residual policy is whether to add the base policy action as an input. Both methods were commonly used in prior residual policy work, some report improvements when the base policy action is not included [45]. We tested both variations on both the book flipping and belt assembly task and obtained mixed results. We report the most successful versions (with base action for the book flipping task, without base action for the belt assembly task) in our result section.
Table 1: The effect of whether to include base policy action as input to residual policy
w/ base action w/o base action
Book flipping 100 75
Belt assembly 25 70
@@ A.5 Hardware Design
Our kinesthetic correction hardware setup features a tool interface that allows task-specific tool swapping. For the book flipping task, we designed a customized fork-shaped tool that can easily insert under the books and flip them. For the belt assembly task, we used a standard WSG-50 gripper and fin-ray fingers [8]. An interesting future direction is to leverage generative models for automatic manipulator design [41]. Future work can also incorporate other types of force or tactile sensors, such as capacitive F/T sensors [9] and vision-based tactile sensors [44, 28, 3].

### v3

@@ 4.3 Comparisons
  • Take-over-Correction (HG-DAgger) [kelly2019hg] is another common correction strategy where human demonstrators monitor policy execution and take complete control when failures are anticipated. We implement Take-over-Correction on our Compliant Intervention Interface by cleaning up command buffer to the compliance controller and switching stiffness to zero upon correction starts, so the robot policy does not affect the robot during correction. However, as explained in § 3.1, take-over correction introduces an abrupt transition around control authority switching, which may cause distributional discontinuities in the training data.
@@
  • Fintune Policy with KL Regularization: A recent method [fan2023dpok] that stabilizes finetuning training by encouraging the predicted action to be close to the training data distribution.
@@ 4.4 Key Findings
Finding 2: Residual policy allows additional useful modality to be added during correction. [Compliant Residual] policy performs significantly better than other methods without force (45% higher success rate than the best position-only baseline on the book task and 53% higher on the belt task) as it can both take in force feedback that indicates critical task information and predict adequate contact forces to apply. For example, the last stage of the book flipping task requires the robot to firmly push the book against the shelf wall to let it stand on its own. [Compliant Residual] policy predicts large pushing forces at this stage to make the books stand stably with a 100% success rate, while [Residual]’s success rate drops from 70% to 35% (§ A.2). The second stage of the belt assembly task (threading the belt on the large pulley) requires delicate belt height adjustments under ambiguous visual information due to occlusions and the lack of depth. [Compliant Residual] policy learns to move along the pulley to find the slot when the finger touches the top of the pulley.
@@
Finding 3: Smooth On-Policy Delta data enables stable residual policy. [Compliant Residual] policy has 45% higher success rate when trained on [On-Policy Delta] data instead of [Take-over-Correction] data on the book flipping task. Residual policy trained with [Take-over-Correction] data sometimes exhibits large noisy motions that trigger task failures, such as retracting the fingers too early in the book flipping task. On the contrary, residual policy trained with [On-policy Delta] data have much smoother action trajectories and better reflect human’s correction intentions, providing suitable magnitudes of corrections.
@@
Finding 4: Retraining base policy is stable but learns correction behavior slowly. Retraining from scratch with the initial and correction data together leads to policies with stable motions. However, its behavior is less affected by the small amount of correction data compared to the dominant portion of initial data, leading to insignificant improvements over the base policy (1.67% success rate drop on the book task averaged across all data collection methods, 18.8% success rate improve on the belt task, both are much less improvements than [Compliant Residual]).
@@
Finding 5: Finetuning base policy is unstable. Policy finetuning with either correction data has the worst performance across all policy update methods and even underperforms the base policy (30% success rate drop on the book task averaged across all data collection methods, 15.6% drop on the belt task). The finetuned policy predicts unstable and noisy motions, quickly leading to out-of-distribution states, such as inserting too high in the book flipping task and drifting away from the board in the belt assembly task. This is likely due to the distribution mismatch between the base policy training data and correction data, causing training instabilities. Adding KL-regularization effectively reduced the noisy behavior, however, the overall success rate is still lower than other baselines.
@@ 4.5 Ablations
Training frequency and batch size. One important parameter in DAgger is the batch size between policy updates. With a smaller batch size, the policy is updated more frequently, then new correction data can better reflect the most recent policy behavior. However, DAgger with small batch sizes is known to suffer from catastrophic forgetting [kirkpatrick2017overcoming, goodfellow2013empirical] since it finetunes neural networks on data with non-stationary distribution. Common solutions include retraining the residual policy at the end of DAgger using all available correction data collected from all the intermediate residual policies [wu2025robocopilot]. Another way is to rely on the base policy training data as a normalizer [he2025uncertainty]. In this work, we empirically found that larger batch sizes can effectively stabilize residual training. With batch size = 50, the book flipping task reach 100% with one batch, while the belt assembly task performs better with two batches. We compare our method with a smaller batch size on the book flipping task, where we warm up the residual with 20 episodes of initial correction data, then update every ten more episodes for three times.
Figure 7: Effect of Training Frequency and Sample. Batch size=50 leads to more stable training and dense sampling after correction starts achieves better performance on the book flipping task.
Finding: Large-batch DAgger is more suitable for training Compliant Residual Policy. The small-batch training becomes unstable and the demonstrator needs to provide large magnitudes of corrections as the number of iterations increases. During evaluation, the final policy always fails by inserting too high, while our single-batch policy achieves a 100% success rate with the same amount of data and training epochs.
Sampling strategy during training. The start of a human intervention contains critical information of the timing and direction of correction. Accurate delta action predictions right after correction starts are important for reactive corrective behaviors. We investigate three strategies for sampling from online correction data during training: 1. Uniform sample, where the whole episode is sampled uniformly. 2. Denser sample around the start of a human intervention, and 3. denser sample only after the human intervention starts. For 2 and 3, we uniformly increase the sample frequency four times for a fixed period before and/or after intervention starts.
@@ 5 Conclusion and Discussion
In this work, we evaluate practical design choices for DAgger in real-world robot learning, and provide a system, CR-DAgger, to effectively collect human correction data with a Compliant Intervention Interface and improve the base policy with a Compliant Residual Policy. We demonstrate the effectiveness of our designs by comparing them with a variety of alternatives on four contact-rich manipulation tasks.
Limitations and Future Work.
The base policy should have a reasonable success rate for the residual policy to learn effectively. From our experiments, we recommend starting to collect correction data for the residual policy when the base policy has at least $10\%\sim 20\%$ success rate. A future direction is to derive theoretical guidelines for the trade-off between the base and residual improvements.
Throughout this work, we use a MLP as the action head of our Compliant Residual Policy and directly regress the actions. Although it works well in our tasks, it may experience difficulty for tasks that involve more distinctive action multi-modalities. More expressive policy formulations, such as Flow Matching [lipman2022flow, black2024pi_0] might be useful for these tasks.
Our data collection system is based on kinesthetic teaching. Although it provides richer data with higher quality than teleoperation as explained in the paper, it may require more labor during training data collection since the demonstrator needs to grasp the handle on the robot.
@@ 6 Acknowledgment
We would like to thank Chendong Xin for his support on experiments and rebuttal. We would like to also thank Eric Cousineau, Huy Ha, and Benjamin Burchfiel for thoughtful discussions on the proposed method, thank Mandi Zhao, Maximillian Du, Calvin Luo, Mengda Xu, and all REALab members for their suggestions on the experiment setup and the manuscript. This work was supported in part by the NSF Award #2143601, #2037101, and #2132519, Sloan Fellowship, and Toyota Research Institute. We would like to thank Google and TRI for the UR5 robot hardware. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
@@ Appendix A Technical Appendices and Supplementary Material
A.1 Control Implementation Details
A.1.1 Compliance Control
Compliance is a motor property that describes how motion responds to force. For example, a balloon has high compliance, meaning that they deform significantly under external forces. Industrial robots typically have very low compliance, meaning that their motion tracking has little deviation under external forces.
Compliance control refers to the technique that makes a rigid robot behave softly using feedback control. It is a standard technique widely adopted in the manufacturing industry. It lets robots interact with the environment safely without creating huge forces. With compliance, the robot retreats when it experiences external forces, larger force leads to bigger position deviations. When there is no external force, compliance control tracks the target position accurately just like position-based control. Compliance control allows specification of the exact desired compliance profile, typically described by parameters such as stiffness, damping and inertia.
Impedance control and admittance control are two methods to implement compliance control. The users are free to choose the controller that works with their robot. Robots with good position control accuracy (e.g., most industrial robots) can use admittance control. Robots with low gear ratio (e.g. "quasi-direct-drive" robots) can use impedance control. We use admittance control in this work and open-sourced our implementation at https://github.com/yifan-hou/force_control. A good reference for different compliance control schemes is the “Force Control” chapter of the Handbook of Robotics [villani2016force].
A.1.2 Control Architecture
Our software system consists of three independent loops:
  1. 1.
    The base policy loop that runs the diffusion policy. The base policy loop updates at about 1Hz, each time predicts 32 frames of actions corresponding to 3.2s of future robot positions. In this work we have not optimized the implementation for computation speed.
  2. 2.
    The residual policy loop that runs at approximately 50Hz, each time predicts five frames of delta actions corresponding to 0.1s of delta positions. The delta actions are added to the corresponding base policy actions based one time, before being sent to the low-level compliance controller for execution. The loop rate is limited by our current implementation and can be improved if needed.
  3. 3.
    The admittance controller loop that runs at exactly 500Hz. This loop implements 6D Cartesian compliance on the robot. It takes reference positions and forces from the residual policy. When there is zero reference force and no external force, the admittance controller lets the robot track the reference position precisely. When external force exists, the robot will deviate from the reference position like a spring centered on the base policy output position.
Apart from the above controller/policy loops, each hardware (e.g. camera, force-torque sensor) has a standalone driver loop maintaining 1. communication with the hardware, and 2. buffers for action and feedback for this hardware.
A.2 Stage-Wise Success Rate
We report the success rate of the book flipping task into three key stages.
@@ A.5 Hardware Design
Our kinesthetic correction hardware setup features a tool interface that allows task-specific tool swapping. For the book flipping task, we designed a customized fork-shaped tool that can easily insert under the books and flip them. For the belt assembly task, we used a standard WSG-50 gripper and fin-ray fingers [chi2024universal]. An interesting future direction is to leverage generative models for automatic manipulator design [xu2024dynamics]. Future work can also incorporate other types of force or tactile sensors, such as capacitive F/T sensors [choi2025coinft] and vision-based tactile sensors [yuan2017gelsight, liu2023enhancing, bauza2024simple].
A.6 Broader Impact
CR-DAgger  contributes to the field of robotics by improving pretrained real-world manipulation policies with a small amount of human correction data. The proposed Compliant Intervention Interface provides an intuitive and safe way for humans to directly interact with robots and correct the robot policy on the spot. We demonstrate significant policy improvements on two real-world contact-rich manipulation tasks, book flipping and belt assembly, which can lead to useful applications in industry and households.